5. An inclined plane is 4 meters long and 1 meter high. A force of 1200 N is needed to move a 3200 N box.
a. Calculate input work.
b. Calculate output work.
c. Calculate IMA.
d. Calculate AMA.
e. Calculate the efficiency of the machine.
1
Expert's answer
2020-04-21T19:02:47-0400

a)


"W_i=1200(4)=4800\ J"

b)


"W_o=3200(1)\frac{}{}=3200\ J"

c)


"IMA=\frac{4}{1}=4"

d)


"AMA=\frac{3200}{1200}=2.67"

e)


"\eta=\frac{3200}{4800}=0.67"